Calculate the amount of time it would take for a 7000 dollar investment to reach 1400 dollars at a simple interest rate of 4% per annum
Found 2 solutions by tommyt3rd, stanbon:
Answer by tommyt3rd(5050)

please clarify, the final amount is less than the initial amount

:)

Answer by stanbon(75887)

Calculate the amount of time it would take for a 7000 dollar investment to reach 1400 dollars at a simple interest rate of 4% per annum
--------
1400 = 7000(1+0.04t)
1 + 0.04t = 2
0.04t = 1
t = 1/0.04
t = 25 years
===================
Cheers,
Stan H.